OpenStudy (bookworm14):

Help with some simple exponents. Can you show me how to handowork 2^1.5 ?

OpenStudy (bookworm14):

I know the answer by using a calculator, but I have to show how to handwork it.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} &= 2\times \sqrt{2} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (triciaal):

1.5 = 3/2

OpenStudy (triciaal):